when 2.1g of a mixture Of KCLO3 and NaCl was thermally decomposed the mass of residue obtained was 1.62g. determine the mass of O2 liberated and calculate the mass if NaCl impurity in the original mixture


1
Expert's answer
2021-03-05T03:29:38-0500

I. Since NaCl is stable to heat, only KClO3 decomposes.

2KClO3----> 2KCl + 3O2

Mass of mixture= 2.1g

Mass of residue= 1.62g

Mass of KClO3 that decomposed= 2.1g - 1.62g = 0.48g

Molar mass of KClO3= 122.5g/mol

Molar mass of O2= 32g/mol

From the balanced equation,

2(122.5)g of KClO3 yields 3(32)g of O2

0.48g of KCLO3 will yield 3(32)/2(122.5) x 0.48 = 0.19g of O2


II. To find the mass of NaCl impurity, let's find the mass of KCl in the residue.

Molar mass of KCl = 74.5g/mol

From the balanced equation,

122.5g of KClO3 yields 74.5g of KCl

0.48g of KClO3 will yield 74.5/122.5 x 0.48 = 0.29g of KCl

Mass of NaCl impurity= 1.62 - 0.29 = 1.33g